Solve by completing the square.. What value should you add to each side of the equation?
x^2+2x=-6

Answer by TimothyLamb(4379)

x^2 + 2x = -6
x^2 + 2x + 6 = 0
---
the above quadratic equation is in standard form, with a=1, b=2, and c=6
---
to solve the quadratic equation by using the quadratic formula, plug this:
1 2 6
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
this quadratic has complex roots:
---
the two complex roots of the quadratic are:
x = -1 + (2.23606798)i
x = -1 - (2.23606798)i
---
check this result by completing the square:
x^2 + 2x = -6
---
convert to the complete-square form:
a(x + d)^2 + e = 0
where:
d = b/(2a) = 2/2 = 1
e = c - (b^2/(4a)) = 6 - (4/4) = 5
a=1, b=2, c=6
---
a(x + d)^2 + e = 0
(x + d)^2 = -e
(x + 1)^2 = -5
x + 1 = sqrt( -5 )
x + 1 = (2.23606798)i
x + 1 - (2.23606798)i = 0
---
x + (1 - (2.23606798)i) = 0
---
the roots of the above equation are:
x = -1 + (2.23606798)i
x = -1 - (2.23606798)i
---
which match the roots obtained from the quadratic formula method above !
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php